OpenStudy (anonymous):

Solve: 4(x - 1) = 2(x + 2) + 6 Answer 14 8 7 28

OpenStudy (anonymous):

4x - 4 = 2x + 4 + 6 4x - 4 = 2x + 10 2x = 14 x = 7

OpenStudy (anonymous):

Thanks I hate my sisters homework